OpenStudy (anonymous):

Integrate 1/((x-1)(x+2))

OpenStudy (anonymous):

= A / (x-1) + B / (x+2) 1 = A(x+2) + B(x-1) let x = 1, A = 1/3 let x = -2, B = -1/3 integrate 1/3 / (x-1) - 1/3 (x+2) = 1/3 * (ln|x-1|) - 1/3 * (ln|x+2|) + c = ( ln |(x-1)/(x+2)| ) / 3

OpenStudy (anonymous):

amirite?

OpenStudy (anonymous):

Yes , thank you